# Can Large Language Models Help Students Prove Software Correctness? An Experimental Study with Dafny

### v3

@@ 5 RQ2 LLM Interaction
In total, the ChatGPT interface collected 206 messages: 103 student prompts and 103 ChatGPT responses. To explore how students interact with ChatGPT during deductive verification tasks, we first analyze the coded prompts. The supplementary material provides the full codebook, along with category counts and user distributions
@@
Six students relied solely on prompting without modifying the LLM code. Four made minor edits to LLM output, while another four were largely autonomous, using the LLM sparingly and heavily modifying code. Two students started with partial implementations. In terms of emotional engagement, three students showed signs of
Two students started with partial implementations. In terms of emotional engagement, three students showed signs of frustration (e.g., P2 prompting the LLM with “did not help!” after two repeated wrong responses ) during the interaction, while two expressed clear satisfaction (e.g., P6 prompting LLM after a correct response “it was strengthing the invariants, thank you”).
@@
5.1 RQ2.1 What strategies do more successful students employ when solving Dafny problems?
To explore this, we divided participants into three groups based on their performance in the LLM-based exercise. G1 includes the top 5 performers (scores 19 through 20 ), G2 the next 4 (scores 17 through 18 ), and G3 the remaining 5 participants (scores 10 through 16) that had the lowest scores. In Figure 2 we show the interaction metrics for each group. Due to space restrictions, only the codes (from the codebook) mentioned in this paper are detailed in Table 2. For information about all codes, please consult the supplementary materials [^2] Link to Supplementary Material.
@@
redirect Guides the AI or teaches Dafny concepts
@@
promptingOnly User only interacted with AI until solution
mostlyPrompting Mostly used AI, with minimal manual edits
@@
Analysis of classCopy shows that all 5 participants in G1 included class definitions in their prompts, compared to 3 out of 4 in G2 and 4 out of 5 in G3. Providing this context was often essential for the LLM to generate a correct solution. For instance, P14 from G3 consistently failed to include sufficient information about the class and its methods, forcing the LLM to make assumptions about the surrounding code and ultimately leading to incorrect outputs.
@@
User initiative varied noticeably across groups. In G1, only one participant relied solely on prompting (promptingOnly). Remarkably, this participant (P8) solved both subproblems of the Tree problem with a single prompt that included the entire Dafny file followed by the instruction “solve this”. This level of success in a single attempt was not observed in any other group. A similar case occurred with P6, also from G1, who used a slightly different prompt for the same problem. However, for unclear reasons, P6 did not recognize the correctness of the LLM’s response and instead began directing the model to revise the code. These examples illustrate the non-deterministic behavior of LLMs: identical or similar prompts can yield different outcomes, and even when a correct solution is generated, users may need sufficient domain knowledge to recognize and apply it.
@@ 6 RQ3 Trust
We asked participants if they trusted (or not) ChatGPT’s responses during the Dafny exercise and why. Overall, participants expressed mixed attitudes. Of the 14 students, seven trusted the LLM and seven distrusted it (see Figure 3).
@@
Seven students distrusted ChatGPT, and of these, four distrusted it because of syntax errors, as it often produced invalid Dafny code that required manual correction (“responses (…) contain one or two little errors, they are sometimes easy to find (…) but [I] still need to be careful.” P8). Two participants also distrusted the LLM because it was overconfident when wrong, and one participant mentioned directly hallucinated features (“[it] kept insisting on wrong suggestions, sometimes suggested things that do not exist” P2). For one participant, distrust came from prior experiences with ChatGPT rather than the current study, and another said that ChatGPT is not useful for verification tasks.
@@
Students’ levels of trust may not be fully aligned with their performance. One example is P5, who scored zero without ChatGPT and near-perfect with it. Despite this improvement, the participant still distrusted the LLM’s output. While ChatGPT can help students complete tasks more successfully, it does not automatically increase their confidence because if it produces correct code that students do not fully understand, they may remain skeptical, even in the face of strong objective performance. On the other hand, some students appeared to place significant trust in ChatGPT even without much critical evaluation. For instance, participant P3 noted that they were tired, turned to ChatGPT, and it produced code quickly, so they “strongly trusted” it.
@@ 7.1 Implications for CS Pedagogy
Our results suggest that good scores with LLM do not necessarily reflect deeper understanding or trust. We saw cases of participants who improved their performance with the LLM and still distrusted it. So, while LLMs can enable students to progress where they might otherwise stall, it does not automatically help learning or understanding. This disconnect raises important questions about dependency, as students may rely on LLMs as a substitute for their reasoning, consistent with prior work [6, 18]. Our survey data offer preliminary support for this concern. Four of the five students who scored above 10 (out of 20) on the unaided verification problem reported using LLMs “sometimes” or “rarely” outside the study. In contrast, the one student in this group who used LLMs “most of the time” was the worst.
@@
For educators, this highlights the need to explicitly teach productive prompting strategies rather than simply allowing or discouraging LLM use. Structured prompting templates or examples could scaffold the learning process and help students internalize good practices over time. The prompting strategies that proved effective in our study may be transferable to other programming tasks. Our results suggest that educators should encourage students to think critically before using LLMs, avoid overly complex solutions, and treat the LLM as a support tool. Breaking down larger problems into smaller steps also seems an effective strategy when working with LLMs. Future work should explore whether teaching prompt design directly leads to improved learning outcomes and whether these benefits generalize across domains.
@@
For educators, our insights show that they must consider what students and LLMs can do. Some design decisions in our problems contributed to increased LLM resistance (e.g., constrained the solution to our defined class invariants, predefined contracts, and selected ghost variables  —  which are not usable in the implementation). These elements reduced the degree of freedom in the solution space, a known strategy for LLM resistance [9]. Our results revealed that ChatGPT struggled more with specification tasks than with implementation. Educators can design exercises focused on specifications (e.g., writing contracts) as these may be less vulnerable to LLM assistance. This insight also has broader implications beyond education. In industrial contexts, they motivate a shift towards specification quality and completeness. If a company uses an LLM assistant for verification, they should focus the human input on the specification and design of correctness properties rather than the implementation itself.
@@ 7.2 Limitations
As with all empirical studies, there are some limitations to our research. First, our sample size was relatively small (14 participants), reflecting the voluntary participation of students enrolled in a specialized formal methods course. We advise future work to look into replicating our protocol on a larger scale. Some of our results can be affected by factors related to task sequence or prior familiarity with Dafny. To address these concerns, we designed tasks of comparable difficulty, randomized the order of tasks, assigned ChatGPT access, and conducted multiple pilot sessions. Finally, our study focused on interactions with a single LLM, ChatGPT, accessed through a custom web interface. While this allowed us to control the user experience and capture interaction logs, it may not fully represent how students use LLMs in real-world settings. Our findings may not generalize to other LLM platforms.

### v4

@@ 5 RQ2 LLM Interaction
In total, the ChatGPT interface collected 206 messages: 103 student prompts and 103 ChatGPT responses. To explore how students interact with ChatGPT during deductive verification tasks, we first analyze the coded prompts. The supplementary material provides the full codebook, along with category counts and user distributions. This section is divided into two sub-research questions.
5.1 RQ2.1 How do students interact with ChatGPT when solving deductive verification problems?
@@
Six students relied solely on prompting without modifying the LLM-generated code. Four made minor edits to LLM generated code, while another four were largely autonomous, using the LLM sparingly and performing manual modifications. Two students started with partial implementations. In terms of emotional engagement, three students showed signs of frustration (e.g., P2 prompting the LLM with “did not help!” after two repeated wrong responses) during the interaction, while two expressed clear satisfaction (e.g., P6 prompting LLM after a correct response “it was strengthening the invariants, thank you”).
@@
5.2 RQ2.2 What strategies do more successful students employ when solving Dafny problems?
To explore this, we divided participants into three groups based on their performance in the LLM-based exercise. G1 includes the top five performers (scores 19 and 20), G2 the next four (scores 17 and 18), and G3 the remaining five participants (scores 10 through 16) that had the lowest scores. In Figure˜2 we show the interaction metrics for each group. Due to space restrictions, only the codes (from the codebook) mentioned in this paper are detailed in Table˜2. For information about all codes, please consult the supplementary materials [^2] Link to Supplementary Material.
@@
redirect Guides the LLM or teaches Dafny concepts
@@
promptingOnly User only interacted with LLM until solution
mostlyPrompting Mostly used LLM, with minimal manual edits
@@
Analysis of classCopy shows that all five participants in G1 included class definitions in their prompts, compared to three out of four in G2 and four out of five in G3. Providing this context was often essential for the LLM to generate a correct solution. For instance, P14 from G3 consistently failed to include sufficient information about the class and its methods, forcing the LLM to make assumptions about the surrounding code and ultimately leading to incorrect outputs.
@@
User initiative varied noticeably across groups. In G1, only one participant relied solely on prompting (promptingOnly). Remarkably, this participant (P8) solved both subproblems of the Tree problem with a single prompt that included the entire Dafny file followed by the instruction “solve this”. This level of success in a single attempt was not observed in any other group. A similar case occurred with P6, also from G1, who used a slightly different prompt for the same problem. However, P6 did not recognize the correctness of the LLM’s response and instead began directing it to revise the code. These examples illustrate the non-deterministic behavior of LLMs: identical or similar prompts can yield different outcomes, and even when a correct solution is generated, users may need sufficient domain knowledge to recognize and apply it.
@@ 6 RQ3 Trust
We asked participants if they trusted ChatGPT’s responses during the Dafny exercise and why. Overall, participants expressed mixed attitudes. Of the 14 students, seven trusted the LLM and seven distrusted it (see Figure˜3).
@@
Seven students distrusted ChatGPT, and of these, four distrusted it because of syntax errors, as it often produced invalid Dafny code that required manual correction (“responses (…) contain one or two little errors, they are sometimes easy to find (…) but [I] still need to be careful.”, P8). Two participants distrusted the LLM because it was overconfident when wrong. One participant mentioned directly hallucinated features (“[it] kept insisting on wrong suggestions, sometimes suggested things that do not exist”, P2). For one participant, distrust came from prior experiences with ChatGPT rather than the current study, and another said that ChatGPT is not useful for verification.
@@
Students’ levels of trust may not be fully aligned with their performance. One example is P5, who scored zero without ChatGPT but had a near-perfect solution with it. Despite this improvement, the participant still distrusted the LLM’s output. While ChatGPT can help students complete tasks more successfully, it does not automatically increase their confidence because if it produces correct code that students do not fully understand, they may remain skeptical, even in the face of strong objective performance. On the other hand, some students appeared to place significant trust in ChatGPT without much critical evaluation. For instance, participant P3 noted that they were tired, turned to ChatGPT, and it produced code quickly, so they “strongly trusted” it.
@@ 7.1 Implications for CS Pedagogy
Our results suggest that good scores with LLM do not necessarily reflect deeper understanding or trust. We saw cases of participants who improved their performance with the LLM and still distrusted it. So, while LLMs can enable students to progress where they might otherwise stall, it does not automatically help learning or understanding. This disconnect raises important questions about dependency, as students may rely on LLMs as a substitute for their reasoning, consistent with prior work [6, 18]. Our survey data offers preliminary support for this concern. Four of the five students who scored above 10 (out of 20) on the unaided problem reported using LLMs “sometimes” or “rarely” outside the study. In contrast, the one student in this group who used LLMs “most of the time” performed the worst.
@@
For educators, this highlights the need to explicitly teach productive prompting strategies rather than simply allowing or discouraging LLM use. Structured prompting templates, or examples, could scaffold the learning process and help students internalize good practices over time. The prompting strategies that proved effective in our study may be transferable to other programming tasks. Our results suggest that educators should encourage students to think critically before using LLMs, avoid overly complex solutions, and treat the LLM as a support tool. Breaking down larger problems into smaller steps also seems an effective strategy when working with LLMs. Future work should explore whether teaching prompt design directly leads to improved learning outcomes and whether these benefits generalize across domains.
@@
For educators, our insights show that they must consider what students and LLMs can do. Some design decisions in our problems contributed to increased LLM resistance (e.g., constrained the solution to our defined class invariants, predefined contracts, and selected ghost variables — which are not usable in the implementation). These elements reduced the degree of freedom in the solution space, a known strategy for LLM resistance [9].
Our results revealed that ChatGPT struggled more with specification tasks than with implementation. Educators can design exercises focused on specifications (e.g., writing contracts) as these may be less vulnerable to LLM assistance. This insight also has broader implications beyond education. In industrial contexts, they motivate a shift towards specification quality and completeness. If a company uses an LLM assistant for verification, they should focus the human input on the specification and design of correctness properties rather than the implementation itself.
@@ 7.2 Limitations
As with all empirical studies, there are some limitations to our research. First, our sample size was relatively small (14 participants), reflecting the voluntary participation of students enrolled in a specialized formal methods course. We advise future work to look into replicating our protocol on a larger scale. Some of our results can be affected by factors related to task sequence or prior familiarity with Dafny. To address this, we designed tasks of comparable difficulty, randomized the order of tasks, assigned ChatGPT access, and conducted multiple pilot sessions. Finally, our study focused on interactions with a single LLM, ChatGPT, accessed through a custom web interface. This allowed control over the user experience and interaction logs, but it may not fully represent students’ real-world LLM use. Our findings may not generalize to other LLM platforms.
@@ 8 Conclusion
{credits}
8.0.1 Acknowledgements
This work was financed by National Funds through the FCT - Fundação para a Ciência e a Tecnologia, I.P. (Portuguese Foundation for Science and Technology) within the project VeriFixer, with reference 2023.15557.PEX (DOI: 10.54499/2023.15557.PEX), and by funding from FCT under grants 2024.00375.BD, PRT/BD/153739/2021, and PRT/BD/155045/2024.